OpenStudy (anonymous):

Evaluate the integral (tan x + cot x)^2

OpenStudy (anonymous):

(tan x+1/tan x)^2dx {(tan^2+1)/tanx)}^2dx now put tan x=t sec^2x dx=dt or,(1+tan^2x)dx=dt or, dx=dt/(1+t^2). put these value in equation {(t^2+1)/t}^2.dt/(1+t^2) or (t^2+1)/t^2 dt or, 1/t^2 dt +1dt or,-1/t+t+c or -1/tanx +tan x+c or , tanx-cotx+c

OpenStudy (anonymous):

verification.... diff.of tanx-cotx+c=(secx)^2+(cosecx)^2 =1+(tanx)^2+1+(cotx)^2 =(tanx)^2+(cotx)^2+2 tanx cotx =(tan x + cot x)^2

OpenStudy (anonymous):

convert tanx +cotx to \[sinx/cosx+cosx/sinx = (\sin ^{2}x + \cos ^{2}x)/(cosxsinx)= 2/\sin2x\] Now take the integral of \[4/\sin ^{2}2x\] because sinxcosx =sin2x/2

OpenStudy (anonymous):

forget all that